Suppose P(A) = 0.25 and P(B) =0.15.If A and B are independent events but not mutually exclusive, find P(A and B)
Answer by ikleyn(52781)

.

Under given conditions, if P(A) = 0.25, P(B) = 0.15 and A and B are independent,


    P(A and B) = P(A)*P(B) = 0.25*0.15 = 0.0375.      ANSWER


The condition that events A and B are not mutually exclusive is EXCESSIVE.

Solved, answered and explained.